Screws Inc. has asked for assistance in determining whether the manufacturing process is operating correctly. When the process is operating properly, it produces screws whose weights are normally distributed with a population mean of 5 grams and a population standard deviation of 0.1 grams.The manager wishes to know if the recent change to a new materials supplier has lowered the weight of the screws. In a random sample of 16 screws, the sample mean weight is 4.962 grams. Is there evidence to conclude that lower weight screws are now being produced at a significance level of 5%? O A. The p-value is 0.0643, and we reject the null hypothesis. O B. The p-value is 0.0643, and we fail to reject the null hypothesis. C. The p-value is 0.1286, and we accept the alternative hypothesis that ,yes, the screws appear to have a lower weight D. The p-value is 0.1286, and we fail to reject the null hypothesis.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 5

Ha : \mu < 5

\bar x = 4.962

\mu = 5

\sigma = 0.1

n = 16

Test statistic = z

= (\bar x - \mu) / \sigma / \sqrtn

= (4.962 - 5) / 0.1 / \sqrt16

= -1.52

Test statistic = -1.52

This is the right tailed test .

P(z <-1.52) = 0.0643

P-value = 0.0643

\alpha = 5% = 0.05

P-value > \alpha .  Fail to reject the null hypothesis .

B)